OpenStudy (anonymous):

the perimeter of a rectangle is 130cm and its width is 28cm. find the area of the rectangle.. wat does this mean............

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

The length of all sides of the rectangle adds up to 130cm. The given width is 28cm, so 28cm * 2 = 56cm, so two sides of the rectangle add up to 56cm. That leaves us with 74cm (130 - 56) for the other two sides. And 74 / 2 = 37cm. So, the top and bottom of the rectangle are 37cm each, and the left and right sides are 28cm each. To get the area of a rectangle, you multiply the length and width together, so in this case the area is 37*28 which equals 1036cm^2. So the answer = 1036cm^2.

OpenStudy (anonymous):

(130- 2*28)/2 * 28 1036cm^2

OpenStudy (anonymous):

^ That works too. Haha

jhonyy9 (jhonyy9):

P=2(w+l) 130=2w+2l 130=2*28+2l 130=56+2l 130-56=2l 74=2l l=74/2 l=37 -so than width is 28 cm and the lenght is 37 cm area = 28*37 = 1036 cm^2

OpenStudy (lgbasallote):

P = 2L + 2W A = LW first of all find L L = (P-2w)/2 so L = (130-2(28)/2 L = 130 - 56/2 L = 74/2 L = 37 so A = 37 x 28 A = 1036 cm^2